A landscape architect bought bluegrass seed for $6 per pound and Bermuda grass seed for $3 per pound. how much Bermuda grass must be mixed with 100 pounds of bluegrass to obtain a blend that will cost $5 per pound?
Answer by richwmiller(17219)

6*100+3x=5*(100+x)
600+3x=500+5x
100 =2x
50=x